### v2

# Expanding Horizons in Short Text Analysis: Integrating LLMs and VAEs for Enhanced Topic Modeling

Abstract
We introduce the Extract-Refine-Retrieve-Read (ERRR) framework, a novel approach designed to bridge the pre-retrieval information gap in Retrieval-Augmented Generation (RAG) systems through query optimization tailored to meet the specific knowledge requirements of Large Language Models (LLMs). Unlike conventional query optimization techniques used in RAG, the ERRR framework begins by extracting parametric knowledge from LLMs, followed by using a specialized query optimizer for refining these queries. This process ensures the retrieval of only the most pertinent information essential for generating accurate responses. Moreover, to enhance flexibility and reduce computational costs, we propose a trainable scheme for our pipeline that utilizes a smaller, tunable model as the query optimizer, which is refined through knowledge distillation from a larger teacher model. Our evaluations on various question-answering (QA) datasets and with different retrieval systems show that ERRR consistently outperforms existing baselines, proving to be a versatile and cost-effective module for improving the utility and accuracy of RAG systems.
1 Introduction
In recent years, the field of natural language processing has witnessed remarkable advancements recently, largely attributable to the proliferation of Large Language Models (LLMs). These LLMs, characterized by their extensive scale training on vast corpora, have demonstrated unparalleled capability of understanding human text and producing high-quality responses kaplan2020scaling; clark2022unified. Moreover, they showcase exceptional practicality and scalability across a spectrum of downstream NLP tasks, such as conversational response generation, text summarization, or content recommendation, under only few-shot or zero-shot settings wu2023survey. However, despite their potential, a major limitation of these LLMs is their struggle to accurately capture the dynamic nature of information updates. Since LLMs are pre-trained on static corpora of data they face difficulties when dealing with the latest or less commonly known information that falls outside their initial training scope. This limitation leads to generating responses that contain outdated or incorrect information, or in some cases, entirely fictitious content— a problem often termed as "hallucination" lee2018hallucinations.
In response to this pressing challenge, the concept of Retrieval-Augmented Generation (RAG) RAG has emerged as a promising technique to enhance the functionality and reliability of LLMs. Retrieval augmentation addresses the shortcomings of original LLMs by incorporating external knowledge (i.e., non-parametric knowledge) sources through some information retrieval system and leveraging the in-context learning ability of LLM readers to generate more contextually relevant and accurate responses. For instance, in a conversational setting where a user interacts with an LLM like ChatGPT gpt to inquire about the latest news events. ChatGPT, being reliant on static pre-training knowledge, may struggle to offer real-time updates in response to the user query. RAG mitigates this information gap by retrieving pertinent news articles or data related to the query from external databases. By augmenting the original user question with retrieved knowledge, the final LLM reader is more likely to generate a contextually relevant answer.
While previous research has underscored the efficacy of retrieval augmentation in mitigating hallucination, it introduces its own set of challenges. One such critical challenge in the RAG systems is the existence of a gap between the information retrievable using the given retriever using original user input queries and the information required for generating optimal responses ragsurvey. Consider a scenario where an external document collection comprises three distinct passages labelled Passage $A$, $B$, and $C$, each containing unique knowledge components denoted as knowledge $x$, $y$, and $z$, respectively. Notably, all three passages include keywords related to Knowledge $z$, which the user seeks to obtain. However, if the user enters a poorly formulated query, there exists a possibility that Passage $A$ or Passage $B$ may be retrieved instead of the ideal source, Passage $C$, thereby limiting the LLM reader’s capacity to derive an optimal response. This disparity is referred to as the pre-retrieval gap in RAG and poses a fundamental obstacle to achieving optimal text generation outcomes.
In the quest to bridge this pre-retrieval information gap ragsurvey within the existing RAG systems, the Rewrite-Retrieve-Read (RRR) framework rrr introduces a pivotal paradigm shift by incorporating the concept of query rewriting from the realm of traditional information retrieval into the RAG domain. RRR proposes the integration of a specialized Query Rewriter module into the native RAG system to align user queries with both the retriever and the LLM reader, thereby optimizing the retrieval process and narrowing the pre-retrieval gap between user queries and the information sought by the LLM reader, i.e. the LLM that generates the final responses. This query optimization technique has also seen widespread adoption in more recent RAG systems stepback; ragsurvey. However, while RRR demonstrates notable advancements, only the original query is passed to the rewrite module, which focuses more on rephrasing queries or expanding the search scope. This framework falls short in tailoring optimization strategies for queries based on the specific needs of the LLM reader, overlooking the need for retrieving the most relevant knowledge essential for accurate generation.
To this end, we propose Extract-Refine-Retrieve-Read (ERRR), a simple but effective framework designed for retrieval augmentation systems. The ERRR framework is crafted to bridge the pre-retrieval information gap through tailored query optimization and aims to resolve the inherent limitations of RRR by enabling retrieval based on the specific information needs of the LLM reader. Specifically, it initiates by extracting parametric knowledge from LLMs and employs a specialized query optimizer that refines user queries. This refinement either complements or validates the extracted parametric knowledge, ensuring that only essential information is retrieved for generating accurate responses, and minimizing the retrieval of extraneous information that could degrade the quality of the output.
Additionally, given that many advanced LLMs, such as ChatGPT gpt, function as black-box systems accessible only through inference APIs, our study introduces a trainable ERRR scheme to enhance customizability for different types of queries. This scheme incorporates a smaller, tunable language model as the query optimizer, which not only helps to lower the expenses associated with query optimization for parametric knowledge refinement but also allows for more adaptable customization tailored to each knowledge source.
Our methods undergo thorough evaluation on multiple question-answering (QA) datasets (HotpotQA hotpotqa, AmbigNQ ambigqa, PopQA popqa). All frameworks and baselines evaluated are implemented using GPT-3.5-Turbo gpt3 or T5-Large t5 as the query optimizer, with GPT-3.5-Turbo serving as both the final reader and the source of parametric knowledge. Evaluation is conducted on the Brave Web Search Engine  [^1] https://brave.com/search/api/ and a local dense retriever, Dense Passage Retrieval (DPR) DPR, alongside a static open-source Wikipedia corpus. Our evaluations demonstrate that ERRR consistently boosts the performance of retrieval-augmented LLMs across all tested QA datasets and retrieval systems when compared to the RRR framework. Furthermore, the outcomes highlight the adaptability and versatility of ERRR, showcasing its effectiveness across diverse settings and data sources.
In summary, our key contributions are as follows: (i) We propose Extract-Refine-Retrieve-Read (ERRR), a retrieval augmentation framework designed to boost the effectiveness of RAG systems by optimizing queries to meet the specific knowledge needs of LLMs. (ii) We reveal that ERRR has a high degree of adaptability across diverse settings and data sources, as well as across various retrieval systems. (iii) We introduce a trainable scheme for ERRR, which not only achieves higher performance but also effectively reduces computational costs.
2 Related work
2.1 Retrieval-Augmented Generation
The integration of retrieval modules to access relevant contextual knowledge has played a crucial role in enhancing Large Language Models (LLMs) in recent years. Initially designed for early sequence-to-sequence models, the Retrieval-Augmented Generation (RAG) framework proposed by Piktus et al. RAG has gained substantial traction in the era of LLM. This approach has diversified into a broad array of methods, with ongoing efforts aimed at further augmenting its capabilities. Earlier exploration primarily focused on improving key components, such as upgrading to more powerful pre-trained language models like BERT bert as readers or employing advanced dense retrievers for retrieval tasks DPR. These retrievers encode documents and inputs into dense vectors, facilitating retrieval based on the similarity between the input and retrieved passages.
Recent studies have shifted focus beyond merely enhancing the retriever or reader components, emphasizing the refinement of pre-retrieval and post-retrieval processes. To address the pre-retrieval gap—the disparity between the information retrievable from original queries and the knowledge required for optimal responses—GenRead genread replaces the retrieval module with a knowledgeable LLM, thereby narrowing the gap between the user query and retrieval process. It prompts the LLM to generate contextual information for the query, using these generated documents as retrieval results to formulate the final answer. Self-ask selfask proposes an iterative approach using chain-of-thought prompting to generate self-posed questions that refine the response. For the post-retrieval gap—the challenge of creating optimal responses from given information—strategies include document re-ranking or summarization. For instance, PRCA prca trains a contextual adaptor module to summarize retrieved documents with a black-box LLM reader.
Several studies have also proposed significant modifications to the original RAG pipeline, introducing complex systems that include both pre-retrieval and post-retrieval modules ragfusion, and adapting the pipeline into iterative or recursive frameworks react; selfrag. While these advanced systems demonstrate notable performance enhancements, they incur substantial costs and typically require multiple interactions with LLM. In contrast, our work focuses on refining the single-turn RAG framework, introducing a flexible and trainable module adaptable to existing systems.
2.2 Query Optimization for Retrieval Augmentation
Recent research highlights a significant discrepancy between input queries and LLM readers for RAG systems, especially under the current trend of using off-the-shelf web search tools or black-box LLMs that are difficult to customize rrr. Typically, these input queries often originate directly from users or specific datasets, which could be either poorly formulated or adhere to a static query format. To overcome these challenges, an effective approach is to optimize the query in the pre-retrieval phase, thereby improving the quality of retrieved information and response generation. The Rewrite-Retrieve-Read (RRR) framework, for instance, trains a query rewriting module using an LLM to better align retrieval queries with LLM readers rrr that generate the final response, as illustrated in Figure 1. Additionally, RRR introduces a trainable scheme that employs reinforcement learning with Proximal Policy Optimization to fine-tune a small open-source model based on feedback from the LLM reader, achieving improved results. HyDE addresses the demand for accurate information retrieval by creating hypothetical documents and encoding them through unsupervised contrastive learning for efficient retrieval operations hyde. Furthermore, Step-Back Prompting stepback converts original queries into high-level abstract questions, aiding LLMs in generating better responses for complex queries requiring abstract thinking.
While these efforts have markedly improved the performance of original RAG systems by focusing on query optimization, they often overlook the importance of synchronizing queries with the specific knowledge requirements of the LLM reader. Unlike the RRR framework, our approach includes an additional parametric knowledge extraction step to assess the knowledge possessed by the LLM. We then perform retrieval based on optimized queries to refine this parametric knowledge, thereby further enhancing retrieval-augmented LLMs.
3 Methodology
Figure 1: Overview of Extract-Refine-Retrieve-Read (ERRR). Extract-Refine-Retrieve-Read leverages parametric knowledge of LLMs and utilizes a specialized query optimizer to retrieve the knowledge that better aligns with LLM’s needs.
In this section, we elaborate on the details of Extract-Refine-Retrieve-Read (ERRR), a framework for improving the retrieval-augmented LLMs through query optimization for parametric knowledge refinement. Section 3.1 formally defines the central task addressed by ERRR and introduces its key concepts. The design of the framework is discussed in Section 3.2, where we outline a frozen scheme using a black-box LLM reader and standard web search tools. Additionally, Section 3.3 discusses a trainable scheme of the framework.
3.1 Pre-retrieval Information Gap
A task with retrieval augmentation can be formulated as follows. Given an input query $q$, a set of theoretical golden documents $D$ that has the accurate information to answer query $q$, and a ground-truth answer $a$, we denote:
[…]
However, to obtain the document set $D$, practical implementations often employ a retrieval function $R$ which retrieves documents $R(q)$ from an external knowledge base, and thus the output of a retrieval-augmented system is:
[…]
An inherent challenge arises due to the difference in the quality and relevance of documents retrieved by $R$ compared to the ideal documents set $D$:
[…]
The limitation discussed above describes the problem of the pre-retrieval gap in the original RAG pipeline, wherein the set $R(q)$ may not adequately represent the information necessary for generating the true answer $a$. Therefore, the main objective is to develop a query optimization function $f$ that transforms the initial user query $q$ into one or more optimized queries $f(q)$ such that $R(f(q))$ better approximates the ideal document set $D$.
Previous work like RRR rrr has demonstrated the effectiveness of such query optimization functions, albeit without considering the influence of $\theta$. To this end, ERRR introduces a more tailored query optimization function $f^{\prime}$ that utilizes the parametric knowledge $\theta$ to perform the query optimization and retrieve external knowledge that refines $\theta$ and better aligns with its needs. This can be formulated as:
[…]
where
$C=E(q~{}|~{}\theta)$
and $E$ denotes the parametric knowledge extraction function.
3.2 Extract-Refine-Retrieve-Read
Extract-Refine-Retrieve-Read consists of a four-step pipeline: Parametric Knowledge Extraction, Query Optimization for Parametric Knowledge Refinement, Retrieval, and Generation, as depicted in Figure 1. Detailed technical implementation for each step, covering the models, prompting techniques and training setup, is provided in Section 4.3.
Parametric Knowledge Extraction  Previous studies such as GenRead genread and HyDE hyde demonstrate that LLMs may possess substantial parametric knowledge capable of addressing user inquiries, particularly on popular topics. Inspired by the prompting methods outlined in GenRead, our approach involves a direct strategy where we prompt the LLM reader to produce a pseudo-contextual document containing all the background information. We consider these pseudo-contextual documents as a representation of the LLM’s abstracted parametric knowledge. Although these documents may contain inaccuracies, they provide essential contextual information related to the original queries.
Query Optimization  In this step, we employ an LLM as the query optimizer for parametric knowledge refinement. We prompt the query optimizer to produce one or more optimized queries seeking external knowledge that either validates or supplements the existing parametric knowledge, especially focusing on the validation of time-sensitive information.
Retrieval  To illustrate the adaptability of our module across various retrieval systems and data sources, we utilize two types of retrievers: a black-boxed web search tool and a local dense retrieval system, which are then combined with the original query for processing by the LLM reader.
Generation  We employ an LLM reader to generate the final answer using both the retrieved documents and the original query. Our prompting strategy involves straightforward instruction followed by 1-3 few-shot examples for question answering. These examples are consistently used within each dataset but vary across different datasets to maintain control over the task-specific output format from the LLM reader—for instance, the responses are expected to be concise in certain QA tasks, usually only one or a few words.
3.3 Trainable Scheme
Given that many powerful LLMs operate as black-box systems, significant challenges such as high computational costs, customization limitations, copyright issues, and connectivity problems have arisen. To address these issues, alongside the conventional frozen scheme, we propose a trainable scheme for our pipeline. Specifically, we fine-tuned a smaller, trainable model utilizing knowledge distillation from a high-performing teacher LLM, leveraging its broadly trained outputs as a good starting point and learning template, and intensively training student models on a distillation dataset of QA questions and generated responses to learn the intricate nuances of query optimization. This streamlined model is then integrated into our pipeline to fulfill the role of query rewriting, originally handled by a frozen LLM.
{tabularx}
[…]
Table 1: List of Prompts Used.
4 Experiments
4.1 Datasets and Metrics
ERRR is assessed on three open-domain question-answering (QA) datasets: AmbigQA ambigqa, PopQA popqa, and HotpotQA hotpotqa. Each dataset serves to test different capabilities of the ERRR framework. (i) The AmbigNQ dataset is the disambiguated variant of the Natural Questions (NQ) dataset, where ambiguous questions from NQ are refined into specific queries with minimal constraints. Consistent with procedures used in RRR, we evaluated ERRR using the first 1000 samples of the test set. (ii) PopQA features simpler questions that focus on less popular knowledge topics compared to other QA tasks. Due to the high similarity in sample distributions, we assessed only the first 997 samples of the test set. (iii) The HotPotQA dataset contains complex questions that require multi-hop reasoning. We conducted evaluations across the entire test set. Following the metric usage for three datasets, our method is evaluated by exact match score $EM$ and $F_{1}$ score.
4.2 Baselines and Proposed Frameworks
We evaluated 7 baselines and proposed frameworks, as detailed below: (i) Direct: Directly calling GPT-3.5-Turbo to answer questions. (ii) RAG: The classic Retrieval-Augmented Generation framework RAG. The original user queries are used for retrieval and fed directly to the LLM reader to generate output. (iii) ReAct: A modified RAG framework that intertwines the reasoning and acting capabilities of LLMs to create a more cohesive and effective approach react. This framework can iteratively perform reasoning prompts and actions, such as information retrieval, serving as our comparison baseline. (iv) Frozen RRR: Rewrite-Retrieve-Read framework rrr with a frozen configuration. It employs GPT-3.5-Turbo to rewrite the query and retrieve relevant documents based on these rewritten queries. Then the original query and retrieved documents are used for reading. This serves as our baseline for comparison. (v) Trainable RRR: Trainable rewrite-retrieve-read framework, initiating with a supervised fine-tuned T5-large model. It then applies reinforcement learning to better align the retriever and rewriter using Proximal Policy Optimization (PPO). This serves as our baseline for comparison. (vi) Frozen ERRR: Extract-Refine-Retrieve-Read framework with a frozen configuration, as described in Section 3.2. (vii) Trainable ERRR: Trainable Extract-Refine-Retrieve-Read framework, as described in Section 3.3.
These frameworks are evaluated using a web search tool or a local retriever with a static corpus, as described in Section 3.2. Due to resource limitations, some frameworks were not evaluated under the local dense retriever setting.
[…]
RAG 0.473 0.5842 0.425 0.4704 0.329 0.4424
ReAct 0.477 0.5787 0.451 0.4917 0.344 0.4649
Frozen RRR 0.452 0.5577 0.445 0.4904 0.337 0.4567
Trainable RRR 0.460 0.5577 0.389 0.4238 0.337 0.4548
Frozen ERRR 0.4815 0.5823 0.480 0.5256 0.369 0.4941
Trainable ERRR 0.4975 0.5988 0.485 0.5309 0.372 0.4989
[…]
Table 2: The retrieval system in the above methods is Brave Search API. "Frozen" indicates the rewriter or the query optimizer is GPT-3.5-Turbo, while "Trainable" refers to the rewriter or the query optimizer is a supervised fine-tuned T5 model. Trainable RRR is also trained using proximal policy optimization (PPO) following the original paper. ’*’ indicates that it is evaluated on 500 random questions drawn from HotPotQA due to resource limitation.
4.3 Implementation Details
For all baselines, we utilized GPT-3.5-Turbo as the primary LLM and adhered to their implementation from the original paper. Notably, for the Trainable RRR, we employed the supervised fine-tuned T5 model checkpoint as the base model. This checkpoint, open-sourced by the original authors, has been warmed up and fine-tuned on multiple datasets to function as the query rewriter. Then we replicated their reinforcement learning process since we replaced the original search tool with the Brave Search Engine. These trainings were conducted on the first 1000 data points for each dataset evaluated, with The training parameters set as follows: a learning rate of 2e-5, 3 epochs, and a batch size of 8.
For our proposed methods ERRR, in addition to the settings mentioned in Section 3.2, the following sections outline technical details:
Parametric Knowledge Extraction  To perform parametric knowledge extraction, we use the same prompts from the GenRead paper and choose the top prompt that is most likely to produce pseudo-contextual documents. We outline these extraction prompts in Table 1.
Query Optimization  Our specific prompt structure is detailed in Table 1, where demonstration consists of 2 manually crafted examples. These examples are consistently used across all tests and primarily serve as one or few-shot examples for the query optimizer.
Retrieval  For our web search engine, we opt for the Brave Search Engine, which, although it may provide slightly lower quality results compared to major competitors like Google or Bing, offers a significantly more cost-effective API. This search API retrieves website snippets, simulating a typical user experience of entering a query in a search engine, pressing Enter, and reviewing the top results at a glance. For local retrieval, we utilize WikiDPR, a specialized subset of Wikipedia collections tailored for the Dense Passage Retrieval (DPR) model DPR. This database consists of 21 million passages from Dec. 20, 2018, each limited to 100 words, along with their 768-dimensional embedded vectors. The retrieval process involves converting a query into a DPR embedding and finding the top k vectors with the closest L2 distances. For both systems, we retrieve the top 5 results, concatenate them with the original query, and feed them to the LLM reader.
Generation  Although different prompting strategies may influence the performance of the question-answering task, this aspect is not the primary focus of our study, so we adhere to the same answer prompts used in the RRR rrr framework. The prompts we used are detailed in Table 1.
Trainable Scheme  For Trainable ERRR, we employ T5-Large t5, an open-source model with 770 million parameters, as the query optimizer. We fine-tune this student model using knowledge distillation from GPT-3.5-Turbo. The distillation dataset was assembled by selecting questions from training sets of each QA dataset, with GPT-3.5-Turbo generating the responses under identical settings utilized in the frozen scheme. We also devised a short eliciting prompt, "Rewrite better search queries to acquire or validate the knowledge needed for the question:", serving as an instruction prefix to guide T5 to adapt to the task. To ensure optimal task-specific outcomes, separate T5 models were trained with 3 epochs for each QA dataset, with a learning rate of 1e-4 and a batch size of 4.
[…]
Frozen RRR 0.438 0.5373 0.378 0.4517 0.289 0.3926
Trainable RRR 0.414 0.5203 0.365 0.4242 0.282 0.3764
Frozen ERRR 0.448 0.5473 0.419 0.4685 0.337 0.4482
Trainable ERRR 0.4595 0.5577 0.426 0.4694 0.338 0.4499
[…]
Table 3: Evaluations with WikiDPR as local retrievers. The other setting is the same as Table 2. Due to resource limitations, some baselines were not fully evaluated under this setting.
4.4 Result
The experimental results across three datasets and two different retrieval tools are presented in Table 2 and Table 3. The Frozen ERRR framework outperforms all the baseline methods—Direct, Frozen RRR, and Trainable RRR—regardless of the retrieval system. This underscores the effectiveness of addressing the pre-retrieval gap and demonstrates adaptability across different retrieval systems and data sources. Additionally, to our surprise, the Trainable ERRR framework further outperforms all other baselines and even its teacher model across all three datasets. We attribute this superior performance to the distillation process, which not only allows the student model to learn from the teacher model’s insights but also to generalize more effectively. In this case, the student model compresses the teacher’s knowledge, potentially learning to focus on critical features and patterns while discarding less relevant information. This compression likely leads to such performance, enabling the model to adapt more proficiently to specific optimization tasks.
[…]
Latency 148s 140s 202s 270s
[…]
Table 4: The total cost and total latency of each method that is evaluated on 200 randomly drawn data points from HotPotQA.
Nevertheless, the impact of the ERRR framework is more noticeable on web search retrieval systems, as evidenced by the degree of performance enhancement between web search and dense retrievers detailed in Table 2 and Table 3. Moreover, the results reveal that both Frozen RRR and Trainable RRR underperform the direct method in the PopQA and HotPotQA datasets when using a dense retriever. This underperformance is attributed to the low quality of the results retrieved from the static corpus, which only includes Wikipedia passages with limited length and knowledge span. Additionally, the low quality of retrieval contributes to the retrieval of more irrelevant documents, which subsequently distracts the Large Language Model (LLM) from correctly answering questions. However, such distraction does not cause ERRR to underperform baseline methods. This suggests that the query optimizer mainly optimizes queries to align with the LLM’s informational needs, thereby reducing the likelihood of retrieving too many irrelevant passages. This is particularly beneficial when operating with a document collection of sub-optimal quality. A relevant case study can be found in Appendix A.
4.5 Cost and Latency
Given our method’s emphasis on a conventional single-turn pipeline, it demonstrates superior performance in terms of cost and latency when compared to certain advanced and iterative RAG frameworks. To underscore the cost-efficiency and flexibility of our approach, we conducted a comparative analysis with ReAct react and Self-RAG selfrag. This experiment was carried out on 200 randomly selected questions from HotPotQA. The results presented in Table 4 highlight that while still maintaining commendable performance, Frozen ERRR exhibits significantly lower costs, faster processing times, and greater efficiency than other iterative frameworks. Moreover, Trainable ERRR has the potential to further reduce costs, particularly for large datasets, by leveraging an already fine-tuned query optimizer, thereby saving on an additional LLM call to GPT-3.5-Turbo.
5 Conclusion
In this paper, we present Extract-Refine-Retrieve-Read (ERRR) framework for Retrieval-Augmented Generation (RAG) systems. The ERRR framework is designed to optimize queries, aligning them closely with the specific informational needs of large language models (LLMs) to enhance retrieval augmentation effectiveness. Our experimental results demonstrate that our method surpasses both the naive LLM and native query rewriting framework Rewrite-Retrieve-Read on benchmark datasets such as AmbigQA ambigqa, PopQA popqa and HotPotQA hotpotqa, utilizing both web search tools and a dense retriever with local static corpus. It demonstrated ERRR’s remarkable adaptability across a variety of settings, data sources, and retrieval systems. This flexibility ensures that ERRR can be effectively implemented in diverse operational environments, making it a potential and adaptable component for inclusion in more advanced RAG systems. Additionally, we have developed and implemented a trainable scheme for the ERRR framework. This approach is both cost-effective and efficient as it relies on only a fine-tuned T5 model trained on a moderately sized dataset and surpasses the performance of the frozen GPT-3.5-Turbo.
6 Limitation
We acknowledge that we recognize the existence of more sophisticated Retrieval-Augmented Generation (RAG) approaches such as Self-RAGselfrag and CRAGCRAG. These advanced systems typically require iterative invocations of the entire pipeline to refine their answers, resulting in exceptionally high computational demands. Due to computational constraints within our study, we focused solely on scenarios that operate in a single-turn manner, wherein each module is invoked only once.
Additionally, our model does not employ any reinforcement learning techniques to enhance the performance of the supervised fine-tuned model. This decision was driven by resource limitations and observed sub-optimal performance when training with a small portion of the dataset using Proximal Policy Optimization (PPO) schulman2017proximal, which constrained the potential upper limit of our model’s performance.
For the future development of this work, while the ERRR framework addresses the pre-retrieval gap problem, future work could extend to methods that bridge the post-retrieval gap or incorporate ERRR into more advanced and modular RAG systems to further enhance performance in question-answering tasks. Furthermore, exploring new Reinforcement Learning (RL) algorithms to improve the query optimizer’s performance for specialized tasks is also a possible direction for further exploration.
{tabularx}
[…]
RRR rewritten query: "actor comedian "The Office" Stories USA cast"
Output: "Ricky Gervais" (incorrect)
Parametric Knowledge Extraction: Stories USA starred actor and comedian Steve Carell from "The Office." Steve Carell is best known for his role as Michael Scott on the hit TV show "The Office," but he has also appeared in a number of films and other television shows. In Stories USA, Carell plays a man who is struggling to come to terms with his past and find his place in the world. The film is a collection of short stories that explore different aspects of the American experience, and Carell’s performance is praised for its depth and emotional resonance. Overall, Stories USA is a thought-provoking and engaging film that showcases Carell’s talent as both a comedic and dramatic actor.
ERRR rewritten queries: "actor and comedian from "The Office" in Stories USA" "Steve Carell role in Stories USA"
Output: "Steven John Carell" (correct)
[…]
RRR rewritten query: "Pakistani actor writer Islamabad Coke Kahani 2012"
Output: "Ali Abbas"(incorrect)
Parametric Knowledge Extraction: One of the Pakistani actors and writers from Islamabad who helped write for the 2012 Pakistani comedy drama sitcom, "Coke Kahani," is Faisal Rehman. Faisal Rehman is a talented actor, writer, and director in the Pakistani entertainment industry. He has appeared in numerous television dramas and films, showcasing his versatility and acting skills. In addition to his acting career, Faisal Rehman has also written scripts for various television shows, including "Coke Kahani," which was well-received by audiences for its humor and engaging storyline. ERRR rewritten queries: "Pakistani actor and writer from Islamabad who helped write for Coke Kahani" "Faisal Rehman contributions to Coke Kahani"
Output: "Yasir Hussain"(correct)
[…]
Table 5: Case Study from RRR and ERRR framework.
Appendix A Case Study
To explicitly and intuitively demonstrate the effectiveness of the ERRR compared to the RRR framework, we present two examples in Table 5 comparing their rewritten queries and final outputs. In the first example, the original question is Stories USA starred which actor and comedian from "The Office"?. The query rewriter in RRR framework produces a simplified query, actor comedian "The Office" Stories USA cast, which merely reformulates the original question for clearer web searching. In contrast, the ERRR not only answers correctly in the Parametric Knowledge Extraction phase but also generates refined queries such as "actor and comedian from "The Office" in Stories USA" and "Steve Carell role in Stories USA", These queries not only attempt to validate the actor name of the T̈he Officeb̈ut also attempt to validate the name Steve Carell from the parametric knowledge, enabling the retriever to source better results.
In the second example, the rewritten query from RRR, Pakistani actor writer Islamabad Coke Kahani 2012, rewrites into only a few random keywords from the original question, which fails to facilitate a high-quality search. On the other hand, the first rewritten query from ERRR, Pakistani actor and writer from Islamabad who helped write for Coke Kahani, provides a clearer and more comprehensible query for search possibly inspired by the contextual information from the extracted parametric knowledge. The second rewritten query, Faisal Rehman contributions to Coke Kahani aims to verify the name derived from parametric knowledge, specifically Faisal Rehman. Interestingly, even though the name is incorrect, the information retrieved subsequently clarifies that Faisal Rehman is not the correct actor and writer, which effectively rectifies the LLM’s output. Together with the information gathered from the first query, this leads to a correct final answer. This example illustrates that even if the pseudo-contextual document contains inaccuracies, the ERRR framework, by concentrating on the specific needs of the LLM reader, can still retrieve the most useful information for the LLM reader which results in a correct outcome.

### v3

@@
-Expanding Horizons in Short Text Analysis: Integrating LLMs and VAEs for Enhanced Topic Modeling
+Query Optimization for Parametric Knowledge Refinement in Retrieval-Augmented Large Language Models
 Abstract
@@
 Our evaluations on various question-answering (QA) datasets and with different retrieval systems show that ERRR consistently outperforms existing baselines, proving to be a versatile and cost-effective module for improving the utility and accuracy of RAG systems.
+Query Optimization for Parametric Knowledge Refinement in Retrieval-Augmented Large Language Models
+Youan Congthanks: Equal contribution.
+Pritom Saha Akash11footnotemark: 1   Cheng Wang   Kevin Chen-Chuan Chang University of Illinois Urbana-Champaign, USA {youanc2, pakash2, chengw4, kcchang}@illinois.edu
 1 Introduction
-In recent years, the field of natural language processing has witnessed remarkable advancements recently, largely attributable to the proliferation of Large Language Models (LLMs).
-These LLMs, characterized by their extensive scale training on vast corpora, have demonstrated unparalleled capability of understanding human text and producing high-quality responses kaplan2020scaling; clark2022unified.
-Moreover, they showcase exceptional practicality and scalability across a spectrum of downstream NLP tasks, such as conversational response generation, text summarization, or content recommendation, under only few-shot or zero-shot settings wu2023survey.
-However, despite their potential, a major limitation of these LLMs is their struggle to accurately capture the dynamic nature of information updates.
-Since LLMs are pre-trained on static corpora of data they face difficulties when dealing with the latest or less commonly known information that falls outside their initial training scope.
-This limitation leads to generating responses that contain outdated or incorrect information, or in some cases, entirely fictitious content— a problem often termed as "hallucination" lee2018hallucinations.
-In response to this pressing challenge, the concept of Retrieval-Augmented Generation (RAG) RAG has emerged as a promising technique to enhance the functionality and reliability of LLMs.
-Retrieval augmentation addresses the shortcomings of original LLMs by incorporating external knowledge (i.e., non-parametric knowledge) sources through some information retrieval system and leveraging the in-context learning ability of LLM readers to generate more contextually relevant and accurate responses.
-For instance, in a conversational setting where a user interacts with an LLM like ChatGPT gpt to inquire about the latest news events.
-ChatGPT, being reliant on static pre-training knowledge, may struggle to offer real-time updates in response to the user query.
-RAG mitigates this information gap by retrieving pertinent news articles or data related to the query from external databases.
-By augmenting the original user question with retrieved knowledge, the final LLM reader is more likely to generate a contextually relevant answer.
-While previous research has underscored the efficacy of retrieval augmentation in mitigating hallucination, it introduces its own set of challenges.
-One such critical challenge in the RAG systems is the existence of a gap between the information retrievable using the given retriever using original user input queries and the information required for generating optimal responses ragsurvey.
-Consider a scenario where an external document collection comprises three distinct passages labelled Passage $A$, $B$, and $C$, each containing unique knowledge components denoted as knowledge $x$, $y$, and $z$, respectively.
-Notably, all three passages include keywords related to Knowledge $z$, which the user seeks to obtain.
-However, if the user enters a poorly formulated query, there exists a possibility that Passage $A$ or Passage $B$ may be retrieved instead of the ideal source, Passage $C$, thereby limiting the LLM reader’s capacity to derive an optimal response.
-This disparity is referred to as the pre-retrieval gap in RAG and poses a fundamental obstacle to achieving optimal text generation outcomes.
-In the quest to bridge this pre-retrieval information gap ragsurvey within the existing RAG systems, the Rewrite-Retrieve-Read (RRR) framework rrr introduces a pivotal paradigm shift by incorporating the concept of query rewriting from the realm of traditional information retrieval into the RAG domain.
-RRR proposes the integration of a specialized Query Rewriter module into the native RAG system to align user queries with both the retriever and the LLM reader, thereby optimizing the retrieval process and narrowing the pre-retrieval gap between user queries and the information sought by the LLM reader, i.e.
-the LLM that generates the final responses.
-This query optimization technique has also seen widespread adoption in more recent RAG systems stepback; ragsurvey.
-However, while RRR demonstrates notable advancements, only the original query is passed to the rewrite module, which focuses more on rephrasing queries or expanding the search scope.
-This framework falls short in tailoring optimization strategies for queries based on the specific needs of the LLM reader, overlooking the need for retrieving the most relevant knowledge essential for accurate generation.
+The field of natural language processing (NLP) has witnessed transformative advancements in recent years, largely driven by the advent of Large Language Models (LLMs).
+These models, trained on vast corpora, have demonstrated exceptional capabilities in understanding human text and generating high-quality responses Kaplan et al.
+(2020); Clark et al.
+(2022).
+They have also proven practical and scalable for various downstream NLP tasks, such as conversational response generation, text summarization, and content recommendation, even in few-shot or zero-shot settings Wu et al.
+(2023).
+Despite their strengths, a key limitation of LLMs lies in their reliance on static training data, which makes them struggle with dynamic or less commonly known information outside their initial training scope.
+This limitation often leads to outdated, inaccurate, or entirely fabricated responses—a phenomenon commonly referred to as “hallucination” Lee et al.
+(2018).
+To address this issue, Retrieval-Augmented Generation (RAG) Lewis et al.
+(2020) has emerged as a promising approach to enhance the functionality and reliability of LLMs.
+By integrating external knowledge sources through retrieval systems, RAG enables LLMs to augment user queries with relevant, up-to-date information.
+This augmentation allows LLMs to generate more contextually accurate and relevant responses.
+For instance, in a conversational setting where a user queries an LLM like ChatGPT Ouyang et al.
+(2022) for the latest news, RAG retrieves pertinent articles to supplement the static pre-trained knowledge of the model, thereby mitigating the information gap.
+While retrieval augmentation has proven effective in mitigating hallucinations, it introduces its own set of challenges.
+A prominent issue in Retrieval-Augmented Generation (RAG) systems is the pre-retrieval gap—a mismatch between the information retrieved using the original user query and the specific knowledge required to generate optimal responses Gao et al.
+(2024).
+For instance, consider a document collection containing three passages, labeled Passage $A$, $B$, and $C$, each containing unique knowledge components $x$, $y$, and $z$, respectively.
+Although all three passages include keywords associated with Knowledge zz—the user’s intended target—a poorly formulated query may lead to retrieving Passage $A$ or $B$ instead of the ideal Passage $C$.
+This misalignment restricts the LLM reader’s ability to generate accurate responses, making the pre-retrieval gap a critical barrier to achieving optimal text generation in RAG systems.
+To bridge the pre-retrieval gap, the Rewrite-Retrieve-Read (RRR) framework Ma et al.
+(2023) introduced query rewriting as a mechanism to optimize user queries and improve their alignment with retrieval systems.
+However, RRR and similar methods Zheng et al.
+(2024); Gao et al.
+(2024) primarily focus on rephrasing or broadening queries, which helps expand the search scope but fails to address the specific knowledge requirements of the LLM reader.
+Additionally, recent self-prompting techniques Li et al.
+(2022); Wang et al.
+(2023) have explored using chain-of-thought (CoT) prompts and pseudo-QA pairs to enhance LLM reasoning capabilities by eliciting internal parametric knowledge.
+While these approaches effectively improve the internal reasoning and explanation capabilities of LLMs for tasks like multi-hop reasoning and open-domain QA, they lack mechanisms for aligning external retrieval queries with the LLM’s knowledge gaps, making them insufficient for resolving the pre-retrieval gap in Retrieval-Augmented Generation (RAG) systems.
 To this end, we propose Extract-Refine-Retrieve-Read (ERRR), a simple but effective framework designed for retrieval augmentation systems.
 The ERRR framework is crafted to bridge the pre-retrieval information gap through tailored query optimization and aims to resolve the inherent limitations of RRR by enabling retrieval based on the specific information needs of the LLM reader.
-Specifically, it initiates by extracting parametric knowledge from LLMs and employs a specialized query optimizer that refines user queries.
+Specifically, it initiates by extracting parametric knowledge from LLMs and employs a specialized query optimizer which refines user queries.
 This refinement either complements or validates the extracted parametric knowledge, ensuring that only essential information is retrieved for generating accurate responses, and minimizing the retrieval of extraneous information that could degrade the quality of the output.
-Additionally, given that many advanced LLMs, such as ChatGPT gpt, function as black-box systems accessible only through inference APIs, our study introduces a trainable ERRR scheme to enhance customizability for different types of queries.
-This scheme incorporates a smaller, tunable language model as the query optimizer, which not only helps to lower the expenses associated with query optimization for parametric knowledge refinement but also allows for more adaptable customization tailored to each knowledge source.
-Our methods undergo thorough evaluation on multiple question-answering (QA) datasets (HotpotQA hotpotqa, AmbigNQ ambigqa, PopQA popqa).
-All frameworks and baselines evaluated are implemented using GPT-3.5-Turbo gpt3 or T5-Large t5 as the query optimizer, with GPT-3.5-Turbo serving as both the final reader and the source of parametric knowledge.
-Evaluation is conducted on the Brave Web Search Engine  [^1] https://brave.com/search/api/ and a local dense retriever, Dense Passage Retrieval (DPR) DPR, alongside a static open-source Wikipedia corpus.
-Our evaluations demonstrate that ERRR consistently boosts the performance of retrieval-augmented LLMs across all tested QA datasets and retrieval systems when compared to the RRR framework.
-Furthermore, the outcomes highlight the adaptability and versatility of ERRR, showcasing its effectiveness across diverse settings and data sources.
-In summary, our key contributions are as follows: (i) We propose Extract-Refine-Retrieve-Read (ERRR), a retrieval augmentation framework designed to boost the effectiveness of RAG systems by optimizing queries to meet the specific knowledge needs of LLMs.
-(ii) We reveal that ERRR has a high degree of adaptability across diverse settings and data sources, as well as across various retrieval systems.
-(iii) We introduce a trainable scheme for ERRR, which not only achieves higher performance but also effectively reduces computational costs.
+In addition to its innovative query optimization process, ERRR introduces a trainable scheme to enhance efficiency and adaptability.
+Recognizing the constraints posed by black-box systems like ChatGPT Ouyang et al.
+(2022), which are accessible only through inference APIs, ERRR incorporates a smaller, tunable language model as the query optimizer.
+This trainable component reduces computational costs while offering greater flexibility to customize the retrieval process for diverse queries and knowledge sources.
+By combining precision in addressing pre-retrieval gaps with cost-effective adaptability, ERRR provides a robust solution for improving retrieval augmentation in LLM-driven systems.
+We evaluate ERRR on multiple question-answering (QA) datasets, including HotpotQA Yang et al.
+(2018), AmbigNQ Min et al.
+(2020), and PopQA Mallen et al.
+(2022), using both web-based (e.g., Brave Search Engine) and local retrieval systems (e.g., Dense Passage Retrieval Karpukhin et al.
+(2020)).
+Across all tested datasets and retrieval configurations, ERRR consistently outperforms baseline frameworks, such as RRR, in terms of retrieval accuracy and response quality.
+These results highlight ERRR’s versatility and effectiveness in diverse settings.
+In summary, our key contributions are as follows: (i) We propose Extract-Refine-Retrieve-Read (ERRR), a novel framework that optimizes queries to bridge the pre-retrieval gap and enhance RAG systems.
+(ii) We demonstrate ERRR’s adaptability across different datasets, retrieval systems, and settings, establishing its robustness and versatility.
+(iii) We introduce a trainable ERRR scheme that reduces computational costs while maintaining high performance, making it suitable for real-world applications.
 2 Related work
@@
 Initially designed for early sequence-to-sequence models, the Retrieval-Augmented Generation (RAG) framework proposed by Piktus et al.
-RAG has gained substantial traction in the era of LLM.
+Lewis et al.
+(2020) has gained substantial traction in the era of LLM.
 This approach has diversified into a broad array of methods, with ongoing efforts aimed at further augmenting its capabilities.
-Earlier exploration primarily focused on improving key components, such as upgrading to more powerful pre-trained language models like BERT bert as readers or employing advanced dense retrievers for retrieval tasks DPR.
+Earlier exploration primarily focused on improving key components, such as upgrading to more powerful pre-trained language models like BERT Devlin et al.
+(2019) as readers or employing advanced dense retrievers for retrieval tasks Karpukhin et al.
+(2020).
 These retrievers encode documents and inputs into dense vectors, facilitating retrieval based on the similarity between the input and retrieved passages.
 Recent studies have shifted focus beyond merely enhancing the retriever or reader components, emphasizing the refinement of pre-retrieval and post-retrieval processes.
-To address the pre-retrieval gap—the disparity between the information retrievable from original queries and the knowledge required for optimal responses—GenRead genread replaces the retrieval module with a knowledgeable LLM, thereby narrowing the gap between the user query and retrieval process.
+To address the pre-retrieval gap—the disparity between the information retrievable from original queries and the knowledge required for optimal responses—GenRead Yu et al.
+(2023) replaces the retrieval module with a knowledgeable LLM, thereby narrowing the gap between the user query and retrieval process.
 It prompts the LLM to generate contextual information for the query, using these generated documents as retrieval results to formulate the final answer.
-Self-ask selfask proposes an iterative approach using chain-of-thought prompting to generate self-posed questions that refine the response.
+Self-ask Press et al.
+(2023) proposes an iterative approach using chain-of-thought prompting to generate self-posed questions that refine the response.
 For the post-retrieval gap—the challenge of creating optimal responses from given information—strategies include document re-ranking or summarization.
-For instance, PRCA prca trains a contextual adaptor module to summarize retrieved documents with a black-box LLM reader.
-Several studies have also proposed significant modifications to the original RAG pipeline, introducing complex systems that include both pre-retrieval and post-retrieval modules ragfusion, and adapting the pipeline into iterative or recursive frameworks react; selfrag.
+For instance, PRCA Yang et al.
+(2023) trains a contextual adaptor module to summarize retrieved documents with a black-box LLM reader.
+Several studies have also proposed significant modifications to the original RAG pipeline, introducing complex systems that include both pre-retrieval and post-retrieval modules Rackauckas (2024), and adapting the pipeline into iterative or recursive frameworks Yao et al.
+(2022); Asai et al.
+(2023).
 While these advanced systems demonstrate notable performance enhancements, they incur substantial costs and typically require multiple interactions with LLM.
@@
 2.2 Query Optimization for Retrieval Augmentation
-Recent research highlights a significant discrepancy between input queries and LLM readers for RAG systems, especially under the current trend of using off-the-shelf web search tools or black-box LLMs that are difficult to customize rrr.
+Recent research highlights a significant discrepancy between input queries and LLM readers for RAG systems, especially under the current trend of using off-the-shelf web search tools or black-box LLMs that are difficult to customize Ma et al.
+(2023).
 Typically, these input queries often originate directly from users or specific datasets, which could be either poorly formulated or adhere to a static query format.
 To overcome these challenges, an effective approach is to optimize the query in the pre-retrieval phase, thereby improving the quality of retrieved information and response generation.
-The Rewrite-Retrieve-Read (RRR) framework, for instance, trains a query rewriting module using an LLM to better align retrieval queries with LLM readers rrr that generate the final response, as illustrated in Figure 1.
+The Rewrite-Retrieve-Read (RRR) framework, for instance, trains a query rewriting module using an LLM to better align retrieval queries with LLM readers Ma et al.
+(2023) that generate the final response, as illustrated in Figure 1.
 Additionally, RRR introduces a trainable scheme that employs reinforcement learning with Proximal Policy Optimization to fine-tune a small open-source model based on feedback from the LLM reader, achieving improved results.
-HyDE addresses the demand for accurate information retrieval by creating hypothetical documents and encoding them through unsupervised contrastive learning for efficient retrieval operations hyde.
-Furthermore, Step-Back Prompting stepback converts original queries into high-level abstract questions, aiding LLMs in generating better responses for complex queries requiring abstract thinking.
+HyDE addresses the demand for accurate information retrieval by creating hypothetical documents and encoding them through unsupervised contrastive learning for efficient retrieval operations Gao et al.
+(2023).
+Furthermore, Step-Back Prompting Zheng et al.
+(2024) converts original queries into high-level abstract questions, aiding LLMs in generating better responses for complex queries requiring abstract thinking.
 While these efforts have markedly improved the performance of original RAG systems by focusing on query optimization, they often overlook the importance of synchronizing queries with the specific knowledge requirements of the LLM reader.
@@
 Given an input query $q$, a set of theoretical golden documents $D$ that has the accurate information to answer query $q$, and a ground-truth answer $a$, we denote:
-[…]
-[…]
+$\text{LLM}(D,q~|~\theta)=a$ (1)
+where LLM denotes an LLM reader and $\theta$ denotes the parametric knowledge of the LLM.
 However, to obtain the document set $D$, practical implementations often employ a retrieval function $R$ which retrieves documents $R(q)$ from an external knowledge base, and thus the output of a retrieval-augmented system is:
-[…]
+$\text{LLM}(R(q),q~|~\theta)$ (2)
 An inherent challenge arises due to the difference in the quality and relevance of documents retrieved by $R$ compared to the ideal documents set $D$:
-[…]
+$\text{LLM}(R(q),q~|~\theta)\neq\text{LLM}(D,q~|~\theta)$ (3)
 The limitation discussed above describes the problem of the pre-retrieval gap in the original RAG pipeline, wherein the set $R(q)$ may not adequately represent the information necessary for generating the true answer $a$.
 Therefore, the main objective is to develop a query optimization function $f$ that transforms the initial user query $q$ into one or more optimized queries $f(q)$ such that $R(f(q))$ better approximates the ideal document set $D$.
-Previous work like RRR rrr has demonstrated the effectiveness of such query optimization functions, albeit without considering the influence of $\theta$.
+Previous work like RRR Ma et al.
+(2023) has demonstrated the effectiveness of such query optimization functions, albeit without considering the influence of $\theta$.
 To this end, ERRR introduces a more tailored query optimization function $f^{\prime}$ that utilizes the parametric knowledge $\theta$ to perform the query optimization and retrieve external knowledge that refines $\theta$ and better aligns with its needs.
 This can be formulated as:
-[…]
+$\text{LLM}(R(f^{\prime}(C,q)),q~|~\theta)$ (4)
 where
@@
 Detailed technical implementation for each step, covering the models, prompting techniques and training setup, is provided in Section 4.3.
-Parametric Knowledge Extraction  Previous studies such as GenRead genread and HyDE hyde demonstrate that LLMs may possess substantial parametric knowledge capable of addressing user inquiries, particularly on popular topics.
+Parametric Knowledge Extraction  Previous studies such as GenRead Yu et al.
+(2023) and HyDE Gao et al.
+(2023) demonstrate that LLMs may possess substantial parametric knowledge capable of addressing user inquiries, particularly on popular topics.
 Inspired by the prompting methods outlined in GenRead, our approach involves a direct strategy where we prompt the LLM reader to produce a pseudo-contextual document containing all the background information.
@@
 This streamlined model is then integrated into our pipeline to fulfill the role of query rewriting, originally handled by a frozen LLM.
-{tabularx}
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+Direct Prompt
+Answer the question in the following format, end the answer with ’**’. {demonstration} Question: {xx} Answer:
+Reader Prompt for Retrieval Augmentation Generation
+Answer the question in the following format, end the answer with ’**’. {demonstration} Question: {docdoc} {xx} Answer:
+Prompt for RRR Query Rewriter
+Think step by step to answer this question, and provide search engine queries for knowledge that you need.
+Split the queries with ’;’ and end the queries with ’**’.
+{demonstration} Question: {xx} Answer:
+Prompt for Parametric Knowledge Extraction
+Generate a background document from web to answer the given question. {xx}
+Prompt for ERRR Query Optimizer
+Address the following questions based on the contexts provided.
+Identify any missing information or areas requiring validation, especially if time-sensitive data is involved.
+Then, formulate several specific search engine queries to acquire or validate the necessary knowledge.
+Split the queries with ’;’ and end the queries with ’**’.
+{demonstration} Context: {Parametric Knowledge} Question: {xx} Queries:
 Table 1: List of Prompts Used.
@@
 4.1 Datasets and Metrics
-ERRR is assessed on three open-domain question-answering (QA) datasets: AmbigQA ambigqa, PopQA popqa, and HotpotQA hotpotqa.
+ERRR is assessed on three open-domain question-answering (QA) datasets: AmbigQA Min et al.
+(2020), PopQA Mallen et al.
+(2022), and HotpotQA Yang et al.
+(2018).
 Each dataset serves to test different capabilities of the ERRR framework.
@@
 We evaluated 7 baselines and proposed frameworks, as detailed below: (i) Direct: Directly calling GPT-3.5-Turbo to answer questions.
-(ii) RAG: The classic Retrieval-Augmented Generation framework RAG.
+(ii) RAG: The classic Retrieval-Augmented Generation framework Lewis et al.
+(2020).
 The original user queries are used for retrieval and fed directly to the LLM reader to generate output.
-(iii) ReAct: A modified RAG framework that intertwines the reasoning and acting capabilities of LLMs to create a more cohesive and effective approach react.
+(iii) ReAct: A modified RAG framework that intertwines the reasoning and acting capabilities of LLMs to create a more cohesive and effective approach Yao et al.
+(2022).
 This framework can iteratively perform reasoning prompts and actions, such as information retrieval, serving as our comparison baseline.
-(iv) Frozen RRR: Rewrite-Retrieve-Read framework rrr with a frozen configuration.
+(iv) Frozen RRR: Rewrite-Retrieve-Read framework Ma et al.
+(2023) with a frozen configuration.
 It employs GPT-3.5-Turbo to rewrite the query and retrieve relevant documents based on these rewritten queries.
@@
 Due to resource limitations, some frameworks were not evaluated under the local dense retriever setting.
-[…]
-[…]
-[…]
+AmbigQA PopQA HotPotQA
+Methods EM F1 EM F1 EM F1
+Direct 0.391 0.4996 0.392 0.4289 0.311 0.4178
 RAG 0.473 0.5842 0.425 0.4704 0.329 0.4424
@@
 Trainable ERRR 0.4975 0.5988 0.485 0.5309 0.372 0.4989
-[…]
 Table 2: The retrieval system in the above methods is Brave Search API.
@@
 For all baselines, we utilized GPT-3.5-Turbo as the primary LLM and adhered to their implementation from the original paper.
+GPT-3.5 Turbo was chosen for its balance of performance and cost, aligning with our focus on optimizing retrieval-augmented generation systems rather than benchmarking generative models themselves.
+While GPT-4 offers improved capabilities, our emphasis remained on augmenting the model’s utility through query optimization.
 Notably, for the Trainable RRR, we employed the supervised fine-tuned T5 model checkpoint as the base model.
@@
 This search API retrieves website snippets, simulating a typical user experience of entering a query in a search engine, pressing Enter, and reviewing the top results at a glance.
-For local retrieval, we utilize WikiDPR, a specialized subset of Wikipedia collections tailored for the Dense Passage Retrieval (DPR) model DPR.
+For local retrieval, we utilize WikiDPR, a specialized subset of Wikipedia collections tailored for the Dense Passage Retrieval (DPR) model Karpukhin et al.
+(2020).
 This database consists of 21 million passages from Dec.
@@
 For both systems, we retrieve the top 5 results, concatenate them with the original query, and feed them to the LLM reader.
-Generation  Although different prompting strategies may influence the performance of the question-answering task, this aspect is not the primary focus of our study, so we adhere to the same answer prompts used in the RRR rrr framework.
+Generation  Although different prompting strategies may influence the performance of the question-answering task, this aspect is not the primary focus of our study, so we adhere to the same answer prompts used in the RRR Ma et al.
+(2023) framework.
 The prompts we used are detailed in Table 1.
-Trainable Scheme  For Trainable ERRR, we employ T5-Large t5, an open-source model with 770 million parameters, as the query optimizer.
+Trainable Scheme  For Trainable ERRR, we employ T5-Large Raffel et al.
+(2020), an open-source model with 770 million parameters, as the query optimizer.
 We fine-tune this student model using knowledge distillation from GPT-3.5-Turbo.
@@
 To ensure optimal task-specific outcomes, separate T5 models were trained with 3 epochs for each QA dataset, with a learning rate of 1e-4 and a batch size of 4.
-[…]
-[…]
-[…]
+AmbigQA PopQA HotPotQA
+Methods EM F1 EM F1 EM F1
+Direct 0.391 0.4996 0.392 0.4289 0.311 0.4178
 Frozen RRR 0.438 0.5373 0.378 0.4517 0.289 0.3926
@@
 Trainable ERRR 0.4595 0.5577 0.426 0.4694 0.338 0.4499
-[…]
 Table 3: Evaluations with WikiDPR as local retrievers. The other setting is the same as Table 2. Due to resource limitations, some baselines were not fully evaluated under this setting.
 4.4 Result
-The experimental results across three datasets and two different retrieval tools are presented in Table 2 and Table 3.
-The Frozen ERRR framework outperforms all the baseline methods—Direct, Frozen RRR, and Trainable RRR—regardless of the retrieval system.
-This underscores the effectiveness of addressing the pre-retrieval gap and demonstrates adaptability across different retrieval systems and data sources.
-Additionally, to our surprise, the Trainable ERRR framework further outperforms all other baselines and even its teacher model across all three datasets.
-We attribute this superior performance to the distillation process, which not only allows the student model to learn from the teacher model’s insights but also to generalize more effectively.
-In this case, the student model compresses the teacher’s knowledge, potentially learning to focus on critical features and patterns while discarding less relevant information.
-This compression likely leads to such performance, enabling the model to adapt more proficiently to specific optimization tasks.
-[…]
-[…]
+Frozen ERRR Trainable ERRR ReAct Self-RAG
+Cost \$0.62 \$0.53 \$1.05 \$1.65
 Latency 148s 140s 202s 270s
-[…]
 Table 4: The total cost and total latency of each method that is evaluated on 200 randomly drawn data points from HotPotQA.
-Nevertheless, the impact of the ERRR framework is more noticeable on web search retrieval systems, as evidenced by the degree of performance enhancement between web search and dense retrievers detailed in Table 2 and Table 3.
-Moreover, the results reveal that both Frozen RRR and Trainable RRR underperform the direct method in the PopQA and HotPotQA datasets when using a dense retriever.
-This underperformance is attributed to the low quality of the results retrieved from the static corpus, which only includes Wikipedia passages with limited length and knowledge span.
-Additionally, the low quality of retrieval contributes to the retrieval of more irrelevant documents, which subsequently distracts the Large Language Model (LLM) from correctly answering questions.
-However, such distraction does not cause ERRR to underperform baseline methods.
-This suggests that the query optimizer mainly optimizes queries to align with the LLM’s informational needs, thereby reducing the likelihood of retrieving too many irrelevant passages.
-This is particularly beneficial when operating with a document collection of sub-optimal quality.
-A relevant case study can be found in Appendix A.
+The experimental results across three datasets and two retrieval tools are presented in Table 2 and Table 3.
+The Frozen ERRR framework consistently outperforms all baseline methods—Direct, Frozen RRR, and Trainable RRR—regardless of the retrieval system used.
+These results highlight the effectiveness of addressing the pre-retrieval information gap, demonstrating ERRR’s adaptability across diverse retrieval systems and datasets.
+Furthermore, the Trainable ERRR framework achieves even better performance, surpassing all baselines and its teacher model (GPT-3.5 Turbo) across all three datasets.
+We attribute this improvement to the distillation process, which enables the student model (fine-tuned T5) to generalize better by focusing on critical features while filtering out irrelevant information.
+This distilled representation allows the model to adapt more effectively to specific query optimization tasks, potentially compressing and refining the teacher’s insights into a more efficient form.
+The impact of the ERRR framework is more pronounced in web search retrieval systems, as evidenced by the greater performance enhancement observed in Table 2 compared to dense retrievers in Table 3.
+This is likely due to the higher quality and broader knowledge span of web-based retrieval systems compared to the static 2018 Wikipedia corpus used for dense retrieval.
+Notably, the results show that both Frozen RRR and Trainable RRR underperform the Direct method in the PopQA and HotPotQA datasets when using dense retrieval.
+This underperformance can be attributed to the low-quality results retrieved from the outdated and limited corpus, which includes only Wikipedia passages of constrained length and scope.
+These limitations lead to an increased retrieval of irrelevant documents, distracting the Large Language Model (LLM) from answering questions correctly.
+In contrast, ERRR demonstrates resilience under such conditions.
+By optimizing queries to align with the LLM’s informational needs, ERRR reduces the retrieval of irrelevant passages, mitigating distractions caused by lower-quality retrieval.
+This robustness is particularly valuable when operating on suboptimal document collections, as it ensures performance gains even in challenging retrieval scenarios.
+A detailed case study, provided in Appendix A, further illustrates how ERRR generates precise queries that enhance retrieval effectiveness and improve final answers, even when the retrieved content includes inaccuracies.
 4.5 Cost and Latency
 Given our method’s emphasis on a conventional single-turn pipeline, it demonstrates superior performance in terms of cost and latency when compared to certain advanced and iterative RAG frameworks.
-To underscore the cost-efficiency and flexibility of our approach, we conducted a comparative analysis with ReAct react and Self-RAG selfrag.
+To underscore the cost-efficiency and flexibility of our approach, we conducted a comparative analysis with ReAct Yao et al.
+(2022) and Self-RAG Asai et al.
+(2023).
 This experiment was carried out on 200 randomly selected questions from HotPotQA.
@@
 The ERRR framework is designed to optimize queries, aligning them closely with the specific informational needs of large language models (LLMs) to enhance retrieval augmentation effectiveness.
-Our experimental results demonstrate that our method surpasses both the naive LLM and native query rewriting framework Rewrite-Retrieve-Read on benchmark datasets such as AmbigQA ambigqa, PopQA popqa and HotPotQA hotpotqa, utilizing both web search tools and a dense retriever with local static corpus.
+Our experimental results demonstrate that our method surpasses both the naive LLM and native query rewriting framework Rewrite-Retrieve-Read on benchmark datasets such as AmbigQA Min et al.
+(2020), PopQA Mallen et al.
+(2022) and HotPotQA Yang et al.
+(2018), utilizing both web search tools and a dense retriever with local static corpus.
 It demonstrated ERRR’s remarkable adaptability across a variety of settings, data sources, and retrieval systems.
@@
 6 Limitation
-We acknowledge that we recognize the existence of more sophisticated Retrieval-Augmented Generation (RAG) approaches such as Self-RAGselfrag and CRAGCRAG.
+We acknowledge that we recognize the existence of more sophisticated Retrieval-Augmented Generation (RAG) approaches such as Self-RAGAsai et al.
+(2023) and CRAGYan et al.
+(2024).
 These advanced systems typically require iterative invocations of the entire pipeline to refine their answers, resulting in exceptionally high computational demands.
@@
 Additionally, our model does not employ any reinforcement learning techniques to enhance the performance of the supervised fine-tuned model.
-This decision was driven by resource limitations and observed sub-optimal performance when training with a small portion of the dataset using Proximal Policy Optimization (PPO) schulman2017proximal, which constrained the potential upper limit of our model’s performance.
+This decision was driven by resource limitations and observed sub-optimal performance when training with a small portion of the dataset using Proximal Policy Optimization (PPO) Schulman et al.
+(2017), which constrained the potential upper limit of our model’s performance.
 For the future development of this work, while the ERRR framework addresses the pre-retrieval gap problem, future work could extend to methods that bridge the post-retrieval gap or incorporate ERRR into more advanced and modular RAG systems to further enhance performance in question-answering tasks.
 Furthermore, exploring new Reinforcement Learning (RL) algorithms to improve the query optimizer’s performance for specialized tasks is also a possible direction for further exploration.
-{tabularx}
-[…]
-[…]
-RRR rewritten query: "actor comedian "The Office" Stories USA cast"
-Output: "Ricky Gervais" (incorrect)
-Parametric Knowledge Extraction: Stories USA starred actor and comedian Steve Carell from "The Office." Steve Carell is best known for his role as Michael Scott on the hit TV show "The Office," but he has also appeared in a number of films and other television shows.
-In Stories USA, Carell plays a man who is struggling to come to terms with his past and find his place in the world.
-The film is a collection of short stories that explore different aspects of the American experience, and Carell’s performance is praised for its depth and emotional resonance.
-Overall, Stories USA is a thought-provoking and engaging film that showcases Carell’s talent as both a comedic and dramatic actor.
-ERRR rewritten queries: "actor and comedian from "The Office" in Stories USA" "Steve Carell role in Stories USA"
-Output: "Steven John Carell" (correct)
-[…]
-[…]
-RRR rewritten query: "Pakistani actor writer Islamabad Coke Kahani 2012"
-Output: "Ali Abbas"(incorrect)
-Parametric Knowledge Extraction: One of the Pakistani actors and writers from Islamabad who helped write for the 2012 Pakistani comedy drama sitcom, "Coke Kahani," is Faisal Rehman.
-Faisal Rehman is a talented actor, writer, and director in the Pakistani entertainment industry.
-He has appeared in numerous television dramas and films, showcasing his versatility and acting skills.
-In addition to his acting career, Faisal Rehman has also written scripts for various television shows, including "Coke Kahani," which was well-received by audiences for its humor and engaging storyline.
-ERRR rewritten queries: "Pakistani actor and writer from Islamabad who helped write for Coke Kahani" "Faisal Rehman contributions to Coke Kahani"
-Output: "Yasir Hussain"(correct)
-[…]
-Table 5: Case Study from RRR and ERRR framework.
 Appendix A Case Study

### v4

@@
 Query Optimization for Parametric Knowledge Refinement in Retrieval-Augmented Large Language Models
-Youan Congthanks: Equal contribution.
-Pritom Saha Akash11footnotemark: 1   Cheng Wang   Kevin Chen-Chuan Chang University of Illinois Urbana-Champaign, USA {youanc2, pakash2, chengw4, kcchang}@illinois.edu
+Youan Cong Pritom Saha Akash Cheng Wang Kevin Chen-Chuan Chang University of Illinois Urbana-Champaign, USA {youanc2, pakash2, chengw4, kcchang}@illinois.edu
 1 Introduction
@@
 (2023).
-Despite their strengths, a key limitation of LLMs lies in their reliance on static training data, which makes them struggle with dynamic or less commonly known information outside their initial training scope.
+Despite their strengths, a key limitation of LLMs lies in their reliance on static training data, which causes them to struggle with dynamic or less commonly known information outside their initial training scope.
 This limitation often leads to outdated, inaccurate, or entirely fabricated responses—a phenomenon commonly referred to as “hallucination” Lee et al.
@@
 While these approaches effectively improve the internal reasoning and explanation capabilities of LLMs for tasks like multi-hop reasoning and open-domain QA, they lack mechanisms for aligning external retrieval queries with the LLM’s knowledge gaps, making them insufficient for resolving the pre-retrieval gap in Retrieval-Augmented Generation (RAG) systems.
-To this end, we propose Extract-Refine-Retrieve-Read (ERRR), a simple but effective framework designed for retrieval augmentation systems.
+To this end, we propose Extract-Refine-Retrieve-Read (ERRR), a straightforward yet effective framework designed for retrieval augmentation systems.
 The ERRR framework is crafted to bridge the pre-retrieval information gap through tailored query optimization and aims to resolve the inherent limitations of RRR by enabling retrieval based on the specific information needs of the LLM reader.
-Specifically, it initiates by extracting parametric knowledge from LLMs and employs a specialized query optimizer which refines user queries.
-This refinement either complements or validates the extracted parametric knowledge, ensuring that only essential information is retrieved for generating accurate responses, and minimizing the retrieval of extraneous information that could degrade the quality of the output.
+Specifically, it initiates by extracting parametric knowledge from LLMs and employs a specialized query optimizer that refines user queries.
+This refinement either complements or validates the extracted parametric knowledge, ensuring that only essential information is retrieved for generating accurate responses, and minimizing the retrieval of extraneous information that could degrade output quality.
 In addition to its innovative query optimization process, ERRR introduces a trainable scheme to enhance efficiency and adaptability.
@@
 The integration of retrieval modules to access relevant contextual knowledge has played a crucial role in enhancing Large Language Models (LLMs) in recent years.
-Initially designed for early sequence-to-sequence models, the Retrieval-Augmented Generation (RAG) framework proposed by Piktus et al.
+Initially designed for early sequence-to-sequence models, the Retrieval-Augmented Generation (RAG) framework proposed by Lewis et al.
 Lewis et al.
-(2020) has gained substantial traction in the era of LLM.
+(2020) has gained substantial traction in the era of LLMs.
 This approach has diversified into a broad array of methods, with ongoing efforts aimed at further augmenting its capabilities.
@@
 For instance, PRCA Yang et al.
-(2023) trains a contextual adaptor module to summarize retrieved documents with a black-box LLM reader.
+(2023) trains a contextual adapter module to summarize retrieved documents with a black-box LLM reader.
 Several studies have also proposed significant modifications to the original RAG pipeline, introducing complex systems that include both pre-retrieval and post-retrieval modules Rackauckas (2024), and adapting the pipeline into iterative or recursive frameworks Yao et al.
@@
 (2023).
-While these advanced systems demonstrate notable performance enhancements, they incur substantial costs and typically require multiple interactions with LLM.
+While these advanced systems demonstrate notable performance enhancements, they incur substantial costs and typically require multiple interactions with LLMs.
 In contrast, our work focuses on refining the single-turn RAG framework, introducing a flexible and trainable module adaptable to existing systems.
@@
 (2023).
-Typically, these input queries often originate directly from users or specific datasets, which could be either poorly formulated or adhere to a static query format.
+Typically, these input queries originate directly from users or specific datasets, which could be either poorly formulated or adhere to a static query format.
 To overcome these challenges, an effective approach is to optimize the query in the pre-retrieval phase, thereby improving the quality of retrieved information and response generation.
@@
 Figure 1: Overview of Extract-Refine-Retrieve-Read (ERRR).
-Extract-Refine-Retrieve-Read leverages parametric knowledge of LLMs and utilizes a specialized query optimizer to retrieve the knowledge that better aligns with LLM’s needs.
-In this section, we elaborate on the details of Extract-Refine-Retrieve-Read (ERRR), a framework for improving the retrieval-augmented LLMs through query optimization for parametric knowledge refinement.
+ERRR leverages parametric knowledge of LLMs and utilizes a specialized query optimizer to retrieve the knowledge that better aligns with the LLM’s needs.
+In this section, we elaborate on the details of Extract-Refine-Retrieve-Read (ERRR), a framework for improving retrieval-augmented LLMs through query optimization for parametric knowledge refinement.
 Section 3.1 formally defines the central task addressed by ERRR and introduces its key concepts.
@@
 We prompt the query optimizer to produce one or more optimized queries seeking external knowledge that either validates or supplements the existing parametric knowledge, especially focusing on the validation of time-sensitive information.
-Retrieval  To illustrate the adaptability of our module across various retrieval systems and data sources, we utilize two types of retrievers: a black-boxed web search tool and a local dense retrieval system, which are then combined with the original query for processing by the LLM reader.
+Retrieval  To illustrate the adaptability of our module across various retrieval systems and data sources, we utilize two types of retrievers: a black-box web search tool and a local dense retrieval system, which are then combined with the original query for processing by the LLM reader.
 Generation  We employ an LLM reader to generate the final answer using both the retrieved documents and the original query.
-Our prompting strategy involves straightforward instruction followed by 1-3 few-shot examples for question answering.
+Our prompting strategy involves straightforward instructions followed by 1-3 few-shot examples for question answering.
 These examples are consistently used within each dataset but vary across different datasets to maintain control over the task-specific output format from the LLM reader—for instance, the responses are expected to be concise in certain QA tasks, usually only one or a few words.
@@
 To address these issues, alongside the conventional frozen scheme, we propose a trainable scheme for our pipeline.
-Specifically, we fine-tuned a smaller, trainable model utilizing knowledge distillation from a high-performing teacher LLM, leveraging its broadly trained outputs as a good starting point and learning template, and intensively training student models on a distillation dataset of QA questions and generated responses to learn the intricate nuances of query optimization.
+Specifically, we fine-tuned a smaller, trainable model using knowledge distillation from a high-performing teacher LLM.
+We leveraged the teacher’s outputs as a strong learning template, intensively training the student model on a distillation dataset of QA questions and generated responses to learn the intricate nuances of query optimization.
 This streamlined model is then integrated into our pipeline to fulfill the role of query rewriting, originally handled by a frozen LLM.
@@
 Frozen RRR 0.452 0.5577 0.445 0.4904 0.337 0.4567
-Trainable RRR 0.460 0.5577 0.389 0.4238 0.337 0.4548
+Trainable RRR 0.460 0.5777 0.389 0.4238 0.337 0.4548
 Frozen ERRR 0.4815 0.5823 0.480 0.5256 0.369 0.4941
@@
 4.3 Implementation Details
-For all baselines, we utilized GPT-3.5-Turbo as the primary LLM and adhered to their implementation from the original paper.
+For all baselines, we utilized GPT-3.5-Turbo as the primary LLM and followed the implementation details from their respective original papers.
 GPT-3.5 Turbo was chosen for its balance of performance and cost, aligning with our focus on optimizing retrieval-augmented generation systems rather than benchmarking generative models themselves.
@@
 Then we replicated their reinforcement learning process since we replaced the original search tool with the Brave Search Engine.
-These trainings were conducted on the first 1000 data points for each dataset evaluated, with The training parameters set as follows: a learning rate of 2e-5, 3 epochs, and a batch size of 8.
+This training was conducted on the first 1000 data points for each dataset evaluated, with The training parameters set as follows: a learning rate of 2e-5, 3 epochs, and a batch size of 8.
 For our proposed methods ERRR, in addition to the settings mentioned in Section 3.2, the following sections outline technical details:
@@
 Frozen ERRR 0.448 0.5473 0.419 0.4685 0.337 0.4482
-Trainable ERRR 0.4595 0.5577 0.426 0.4694 0.338 0.4499
+Trainable ERRR 0.4595 0.5777 0.426 0.4694 0.338 0.4499
 Table 3: Evaluations with WikiDPR as local retrievers. The other setting is the same as Table 2. Due to resource limitations, some baselines were not fully evaluated under this setting.
@@
 This underperformance can be attributed to the low-quality results retrieved from the outdated and limited corpus, which includes only Wikipedia passages of constrained length and scope.
-These limitations lead to an increased retrieval of irrelevant documents, distracting the Large Language Model (LLM) from answering questions correctly.
+These limitations lead to an increased retrieval of irrelevant documents, hindering the LLM from answering questions correctly.
 In contrast, ERRR demonstrates resilience under such conditions.
@@
 The results presented in Table 4 highlight that while still maintaining commendable performance, Frozen ERRR exhibits significantly lower costs, faster processing times, and greater efficiency than other iterative frameworks.
-Moreover, Trainable ERRR has the potential to further reduce costs, particularly for large datasets, by leveraging an already fine-tuned query optimizer, thereby saving on an additional LLM call to GPT-3.5-Turbo.
+Moreover, Trainable ERRR has the potential to further reduce costs, particularly for large datasets, by leveraging an already fine-tuned query optimizer, thereby eliminating an additional LLM call to GPT-3.5-Turbo.
 5 Conclusion
@@
 (2018), utilizing both web search tools and a dense retriever with local static corpus.
-It demonstrated ERRR’s remarkable adaptability across a variety of settings, data sources, and retrieval systems.
+These results demonstrated ERRR’s remarkable adaptability across a variety of settings, data sources, and retrieval systems.
 This flexibility ensures that ERRR can be effectively implemented in diverse operational environments, making it a potential and adaptable component for inclusion in more advanced RAG systems.
@@
 6 Limitation
-We acknowledge that we recognize the existence of more sophisticated Retrieval-Augmented Generation (RAG) approaches such as Self-RAGAsai et al.
+We acknowledge the existence of more sophisticated Retrieval-Augmented Generation (RAG) systems such as Self-RAG Asai et al.
 (2023) and CRAG Yan et al.
@@
 These advanced systems typically require iterative invocations of the entire pipeline to refine their answers, resulting in exceptionally high computational demands.
-Due to computational constraints within our study, we focused solely on scenarios that operate in a single-turn manner, wherein each module is invoked only once.
-Additionally, our model does not employ any reinforcement learning techniques to enhance the performance of the supervised fine-tuned model.
-This decision was driven by resource limitations and observed sub-optimal performance when training with a small portion of the dataset using Proximal Policy Optimization (PPO) Schulman et al.
-(2017), which constrained the potential upper limit of our model’s performance.
-For the future development of this work, while the ERRR framework addresses the pre-retrieval gap problem, future work could extend to methods that bridge the post-retrieval gap or incorporate ERRR into more advanced and modular RAG systems to further enhance performance in question-answering tasks.
-Furthermore, exploring new Reinforcement Learning (RL) algorithms to improve the query optimizer’s performance for specialized tasks is also a possible direction for further exploration.
+Due to these computational constraints, our study focused solely on scenarios that operate in a single-turn manner, wherein each module is invoked only once.
+Additionally, our trainable query optimizer does not employ any reinforcement learning (RL) techniques to further refine its alignment with the reader, a decision driven by resource constraints and observed performance degradation when training on a small subset of the dataset using Proximal Policy Optimization (PPO) Schulman et al.
+(2017).
+The effectiveness of the ERRR framework is also inherently tied to the capabilities of the underlying LLMs used for knowledge extraction, query optimization, and final generation.
+The quality of the extracted knowledge and the refined queries are dependent on the base model’s own internal knowledge and reasoning abilities.
+Consequently, the framework’s overall performance may vary significantly when implemented with different or future LLMs.
+These limitations point toward several directions for future work.
+Beyond addressing the pre-retrieval gap, future research could also tackle post-retrieval challenges by exploring methods to better re-rank, filter, or synthesize retrieved documents before they are passed to the reader.
+Furthermore, ERRR could be incorporated as a modular component within more advanced, iterative RAG systems to create a more holistic and powerful RAG pipeline.
+Finally, exploring novel RL algorithms tailored for this framework could lead to further improvements for the trainable optimizer, allowing it to better realize its potential for adaptation.
+7 Acknowledgement
+This material is based upon work supported by the National Science Foundation IIS 16-19302 and IIS 16-33755, Zhejiang University ZJU Research 083650, IBM-Illinois Center for Cognitive Computing Systems Research (C3SR) and IBM-Illinois Discovery Accelerator Institute (IIDAI), grants from eBay and Microsoft Azure, UIUC OVCR CCIL Planning Grant 434S34, UIUC CSBS Small Grant 434C8U, and UIUC New Frontiers Initiative.
+Any opinions, findings, conclusions, or recommendations expressed in this publication are those of the author(s) and do not necessarily reflect the views of the funding agencies.
 Appendix A Case Study
-To explicitly and intuitively demonstrate the effectiveness of the ERRR compared to the RRR framework, we present two examples in Table 5 comparing their rewritten queries and final outputs.
+To explicitly and intuitively demonstrate the effectiveness of the ERRR compared to the RRR framework, we present two examples in Table LABEL:case_study comparing their rewritten queries and final outputs.
 In the first example, the original question is Stories USA starred which actor and comedian from "The Office"?.
@@
 In contrast, the ERRR not only answers correctly in the Parametric Knowledge Extraction phase but also generates refined queries such as "actor and comedian from "The Office" in Stories USA" and "Steve Carell role in Stories USA", These queries not only attempt to validate the actor name of the T̈he Officeb̈ut also attempt to validate the name Steve Carell from the parametric knowledge, enabling the retriever to source better results.
+Example 1
+Question Stories USA starred which actor and comedian from "The Office"?
+Answer Steven John Carell
+RRR
+Rewritten Query(s) actor comedian "The Office" Stories USA cast
+Output Ricky Gervais (incorrect)
+ERRR
+Parametric Knowledge The model’s internal knowledge correctly identifies Steve Carell as the star of "Stories USA" and an actor from "The Office," providing additional (though unverified) details about his role and career.
+Rewritten Query(s) actor and comedian from "The Office" in Stories USA, Steve Carell role in Stories USA
+Output Steven John Carell (correct)
+Example 2
+Question What Pakistani actor and writer from Islamabad helped write for the 2012 Pakistani comedy drama sitcom, "Coke Kahani"?
+Answer Yasir Hussain
+RRR
+Rewritten Query(s) Pakistani actor writer Islamabad Coke Kahani 2012
+Output Ali Abbas (incorrect)
+ERRR
+Parametric Knowledge The model incorrectly identifies Faisal Rehman as a writer for "Coke Kahani," but correctly notes he is a Pakistani actor and writer.
+This flawed parametric knowledge still provides a useful starting point for query refinement.
+Rewritten Query(s) Pakistani actor and writer from Islamabad who helped write for Coke Kahani, Faisal Rehman contributions to Coke Kahani
+Output Yasir Hussain (correct)
+Table 5: A case study comparing the RRR and ERRR frameworks.
+For each example, we present the query and output from the baseline RRR, followed by the extracted parametric knowledge, refined queries, and final output from ERRR.
 In the second example, the rewritten query from RRR, Pakistani actor writer Islamabad Coke Kahani 2012, rewrites into only a few random keywords from the original question, which fails to facilitate a high-quality search.
@@
 Together with the information gathered from the first query, this leads to a correct final answer.
-This example illustrates that even if the pseudo-contextual document contains inaccuracies, the ERRR framework, by concentrating on the specific needs of the LLM reader, can still retrieve the most useful information for the LLM reader which results in a correct outcome.
+This example illustrates that even if the pseudo-contextual document contains inaccuracies, the ERRR framework, by concentrating on the specific needs of the LLM reader, can still retrieve the useful information, resulting in a correct outcome.